# Schema-based active inference supports rapid generalization of experience and frontal cortical coding of abstract structure

### v1

@@ 4 Methods
In this Section, we first briefly review the functioning of active inference and then we illustrate the structure of the novel schema-based (S-HAI) agent.
@@ 4.2.3 Clone-structured causal graphs
We also implement a mechanism to learn the CSCGs online in Section 2.4. In contrast to standard parameter learning in active inference, where the belief over the state is filtered as actions are executed and observations come in, clone-graphs smooth the belief over states, and propagate disambiguating information back into other “cloned” states. This provides a better estimate over the individual states, which can then be used to update the Dirichlet distribution over the transition parameters. For learning the schemas (i.e., the level 2 transitions), we use a sliding window of 10 observations and update the parameters as each observation comes in.
@@ Acknowledgments
This research received funding from the European Research Council under the Grant Agreement No. 820213 (ThinkAhead), the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union – NextGenerationEU (Project IR0000011, CUP B51E22000150006, ‘EBRAINS-Italy’; Project PE0000013, ‘FAIR’; Project PE0000006, ‘MNESYS’), and the Ministry of University and Research, PRIN PNRR P20224FESY and PRIN 20229Z7M8N. The GEFORCE Quadro RTX6000 and Titan GPU cards used for this research were donated by the NVIDIA Corporation. We used a Generative AI model to correct typographical errors and edit language for clarity.

### v2

@@ 4 Methods
In this section, we first briefly review the functioning of active inference, and then we illustrate the structure of the novel schema-based (S-HAI) agent.
@@ 4.2.3 Clone-structured causal graphs
We also implement a mechanism to learn the CSCGs online in Section 2.4. In contrast to standard parameter learning in active inference, where the belief over the state is filtered as actions are executed and observations come in, clone-graphs smooth the belief over states, and propagate disambiguating information back into other “cloned” states. This provides a better estimate of the individual states, which can then be used to update the Dirichlet distribution over the transition parameters. For learning the schemas (i.e., the Level 2 transitions), we use a sliding window of 10 observations and update the parameters as each observation comes in.
@@ Acknowledgments
We thank Mohamady El-Gaby and Adam Harris for useful feedback and discussions. This research received funding from the European Research Council under the Grant Agreement No. 820213 (ThinkAhead), the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union – NextGenerationEU (Project IR0000011, CUP B51E22000150006, ‘EBRAINS-Italy’; Project PE0000013, ‘FAIR’; Project PE0000006, ‘MNESYS’), and the Ministry of University and Research, PRIN PNRR P20224FESY and PRIN 20229Z7M8N. The GEFORCE Quadro RTX6000 and Titan GPU cards used for this research were donated by the NVIDIA Corporation. We used a Generative AI model to correct typographical errors and edit language for clarity.